OpenStudy (andrewthecookie):

You have $59.95 in your wallet and want to buy some new CDs. If the CDs are 11.99 dollars each, what number of CDs, x, can you buy? Write and solve an inequality.

OpenStudy (andrewthecookie):

@AriPotta

OpenStudy (andrewthecookie):

@wio @perl @Zarkon @Xaze @calculusfunctions @M4thM1nd @quickstudent

Nnesha (nnesha):

total money you have is 59.95

OpenStudy (anonymous):

\[0 \le x \le 59.95/11.99\] \[0 \le x \le 5\]

Nnesha (nnesha):

and